OpenStudy (anonymous):

The graph below shows a system of equations: Draw a line labeled 2 y equals minus x minus 1 by joining the ordered pairs negative 3,1 and 3, negative 2. Draw a line labeled 2 y equals x plus 5 by joining the ordered pairs negative 3, 1 and 3, 4 The x-coordinate of the solution to the system of equations is _____ Numerical Answers Expected!

OpenStudy (anonymous):

@lacrosseplayer22

OpenStudy (anonymous):

Hey I closed the question, can you just answer this?

Elsa213 (elsa213):

@camerondoherty

OpenStudy (camerondoherty):

I need the graph...

OpenStudy (anonymous):

Im on FLVS, how do I send over the graph?

OpenStudy (camerondoherty):

save image and attach the file

Elsa213 (elsa213):

@BoyWhoWantsAnA

OpenStudy (anonymous):

I saved it and I put it on here, but it never loaded...

OpenStudy (anonymous):

are these the 2 lines? 2y=-x-1 2y=x+5 if so....change the equations to slope intercept form. so the frist one would be y=-1/2x-1/2 and the second would be y= 1/2x+5/2  is that how the graph looks like. sorry its really messy and all